OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 2) and (1, 5)? y − 5 = 3(x − 1) y − 2 = −3(x − 0) y + 2 = 3(x + 1) y + 5 = −3(x + 1) @amistre64 @ganeshie8 @iambatman Can anyone plz help me?

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\] point slope form, so first use the slope equation then plug in for (y_1,x_1) slope (m) equation \[ m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

(x_1,y_1) haha

OpenStudy (anonymous):

but in the answer choices the it's written like this y − 5 = 3(x − 1) and the rest is similar but different numbers. the answer can't be (x1,y1) @iambatman

OpenStudy (anonymous):

Yeah so fine the slope first and plug in (x_1,y_1)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So you can plug in (0,2) for your x_1, y_1 but you still have to find slope m.

OpenStudy (anonymous):

i got m= 3

OpenStudy (anonymous):

You can do the same with other coordinates as well, so you have two answers

OpenStudy (amistre64):

by point slope definition, we can narrow it to the other 2, and pick the 3 slope

OpenStudy (anonymous):

yes A and C

OpenStudy (anonymous):

Why c?

OpenStudy (anonymous):

i meant the 2 answers that we know r right because of the slope.

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

what is it?